### v1

# Rydberg Atomic Quantum Receivers for Classical Wireless Communication and Sensing

Abstract
The Rydberg atomic quantum receiver (RAQR) is an emerging quantum precision sensing platform designed for receiving radio frequency (RF) signals. It relies on creation of Rydberg atoms from normal atoms by exciting one or more electrons to a very high energy level, which in turn makes the atom sensitive to RF signals. The RAQR realizes RF-to-optical conversion based on light-atom interaction relying on the so called electromagnetically induced transparency (EIT) and Aulter-Townes splitting (ATS), so that the desired RF signal can be read out optically. The large dipole moments of Rydberg atoms associated with rich choices of Rydberg states and various modulation schemes facilitate an ultra-high sensitivity ($\sim$ nV/cm/$\sqrt{\text{Hz}}$) and an ultra-broadband tunability (near direct-current to Terahertz). RAQRs also exhibit compelling scalability and lend themselves to the construction of innovative, compact receivers. Initial experimental studies have demonstrated their capabilities in classical wireless communications and sensing. To fully harness their potential in a wide variety of applications, we commence by outlining the underlying fundamentals of Rydberg atoms, followed by the principles, structures, and theories of RAQRs. Finally, we conceive Rydberg atomic quantum single-input single-output (RAQ-SISO) and multiple-input multiple-output (RAQ-MIMO) schemes for facilitating the integration of RAQRs with classical wireless systems, and conclude with a set of potent research directions.
I Introduction
Next-generation (NG) wireless systems are expected to support a variety of functionalities and a multitude of applications. Integrated communication and sensing functionalities of NG wireless require particularly high sensitivity and data rates, while supporting large-scale connectivity and ultra-low latency. Therefore, the receivers should possess broadband tunability and wideband processing capability. Based upon advanced antennas and integrated circuit technologies, cutting-edge radio frequency (RF) receivers support RF signal reception by well-calibrated antennas, filters, amplifiers, and mixers [1]. The sensitivity is limited by the combination of the receiver’s bandwidth, noise figure, and the minimum signal-to-noise ratio (SNR) required to demodulate the desired signal. Given a specific sensitivity, strong RF reception capability can be attained by using multi-antenna based multiple-input multiple-output (MIMO) schemes and sophisticated signal processing methods to combat channel fading and noise contamination. To receive wideband RF signals at different carrier frequencies, both broadband tunability and wideband processing capability may be realized by stacking several integrated circuit branches along with extra metallic components, where each branch is responsible for a specific frequency band. Such RF systems are however band-limited and bulky, power-thirsty, and suffer from numerous design challenges, especially for large-scale, high-frequency receivers.
The Rydberg atomic quantum receiver (RAQR) concept emerges as a radical solution to these RF reception challenges [2, 3]. It relies on the creation of Rydberg states (excitation), which have remarkably high dipole moment and hence are eminently suitable for sensing the impinging RF signals. The response to an RF signal in RAQR is detected optically via change in the atomic spectroscopy signal. The RF field’s amplitude measurements are directly linked to the International System of Units (SI) and they set absolute (atomic) standards for the RF electric fields from a metrology standpoint. Additionally, they are capable of directly down-converting RF signals, spanning from kilohertz (kHz) to Terahertz (THz), to the baseband without using any mixers, hence significantly simplifying the receiver’s structure and yielding a compact form-factor relative to conventional RF receivers.
To fully unlock the potential of RAQRs in communications and sensing, we offer a panoramic overview of RAQRs. We commence by highlighting the fundamentals of Rydberg atoms and proceed to powerful receivers, including their principles, structures, theories, impairments, and their critical comparisons to RF receivers. We then summarize their capabilities and distinctive features, surveying typical studies from diverse research communities. Furthermore, we conceive an Rydberg atomic quantum single-input single-output (RAQ-SISO) system and a potential RAQ-MIMO scheme to facilitate the integration of RAQRs with existing classical wireless systems. We finally forecast several promising future directions in this very first article on RAQRs in wireless communication and sensing from an engineering perspective.
Figure 1: (a) Energy level diagram showing sub-levels of an atom. (b) Scaling of the relevant properties of Rydberg states with the alkali atom principal number ($n$). (c) Radial probability distribution of Rydberg states of Cs atoms for $n=10,90$. (d) Electron probability distribution $|\psi_{nlm}|^{2}$ with $n=70$, $l=50$, $m=40$, for hydrogen-like atoms.
II Rydberg Atomic Quantum Receiver
RARQs exploit the properties of Rydberg atoms, hence we commence by their brief overview, followed by their principles, structures and theories. Finally, we present the impairments of RAQRs, and compare them to classical RF receivers.
II-A Fundamentals of Rydberg Atoms
II-A1 Atomic Structure [4]
Quantum mechanics provides a comprehensive description of the structure, properties and behavior of atoms. Each atom has a nucleus at its core and there are electrons that orbit around the nucleus. The energy of electrons in these orbits is quantized, i.e., they occupy only certain discrete energy-levels and cannot be traced in between these levels. However, electrons can transition from one energy level to another by absorbing or emitting a photon with energy equal to the difference between the energy levels. Electrons have a wave-particle dual nature. In the quantum mechanical description, the state of an electron is described by a wavefunction $\psi$, which contains all information about the electron’s position, energy, spin, orientation, etc. The square of the wavefunction, $|\psi|^{2}$ gives the probability density (i.e. probability per unit volume) of finding the electron in a certain region of space around the nucleus. These regions are termed as orbitals. The unique quantum state of an electron is described by four quantum numbers:
  • Principal quantum number $\bm{n}$ (positive integer $n$): It indicates the energy level or shell which the electron occupies. A higher $n$ implies that the electron resides further away from the nucleus.
  • Azimuthal or orbital angular momentum quantum number $\bm{l}$ (non-negative integer $l\leq n-1$): It accounts for the sub-shell (sub-level) that the electron occupies, with each sub-shell having unique characteristic shapes, as described in [4]. The letters $S,P,D,F,G,H,I,...$ are assigned to $l=0,1,2,3,4,5,6,...$ respectively.
  • Magnetic quantum number $\bm{m}$ (integer $-l\leq m\leq l$): It indicates the specific orbital that an electron occupies, which corresponds to a certain spatial orientation.
  • Spin quantum number $\bm{s}$ ($s=\pm\frac{1}{2}$): It denotes the orientation of an electron’s intrinsic spin. Following Pauli’s exclusion principle, two electrons in an identical orbital must have opposite spin orientations.
The energy level structure with orbitals up to $6P$ is shown in Fig. 1(a). If mediated by photons of electromagnetic (EM) field, the transition of an electron from one quantum state to another is called an electric dipole transition. Based on the conservation of angular momentum and parity [^1] Parity refers to the symmetry of a wavefunction under spatial inversion., only those (electric dipole) transitions are allowed which satisfy the selection rules: ① $\Delta l=\pm 1$, ② $\Delta m=0,\pm 1$, and ③ the initial and final states must have opposite parities.
II-A2 Properties of Rydberg State
Alkali atoms have single valence electrons in their ground state in the outermost orbital. This represents a relatively loose bound and therefore it can be excited to a high energy level associated with a high principal quantum number, of say $n>10$ for atomic systems. In such a situation, the atom is a Rydberg atom and the corresponding atomic state is a Rydberg state. The Rydberg states of any alkali atom and a hydrogen atom have essentially similar structures and properties since in both cases there is a single electron occupying a high-$n$ state and having an effective core of unit positive charge at the center. The only difference lies in the finite size of the positively charged core, which is larger for alkali atoms than for hydrogen atoms. This leads to a small correction (via quantum the defect parameter $\delta_{l}$) to the energies for non-hydrogenic atoms in particular for low $l$ states. But again, one can say that the general properties of all Rydberg atoms are quite similar. The scaling of the relevant properties of Rydberg states with $n$ is shown in Fig. 1(b). The $n$-dependence arises from the Rydberg atom wavefunction characteristics. For RAQRs, the typical choice of alkali atoms falls on either Caesium (Cs) or Rubidium (Rb) due to the availability of commercial laser systems for the transition wavelengths of these species.
The Rydberg atoms are large, since the average distance from the nucleus $\langle r\rangle$ scales with $n^{2}$, resulting in large dipole matrix elements. This point is illustrated in Fig. 1(c), where the radial probability distribution $|rR_{nl}(r)|^{2}$ for state $nS$ of Cs atoms is plotted against $r$ (distance from nucleus) for increasing $n$. The atomic size increases dramatically with $n$ and its value approaches 1 $\mu m$ for $n=100$, while the ground-state atom is just 0.1 nm in size. The spatial extent and particular shape of the electron probability distribution $|\psi_{nlm}|^{2}$, for $n=70$, $l=50$, $m=40$, is captured in Fig. 1(d) for hydrogen atom. The Rydberg-atom wavefunction (with same quantum numbers) for any alkali atom would closely resemble to this electronic distribution for such high $l$ and $m$ states.
Figure 2: Stark map of Cs atoms, in the vicinity of the $30D$ state, showing the energy shifts experienced by different states in the presence of electric field. The energy level $30D$ in the absence of electric field is marked by red dashed line
II-A3 Rydberg Atoms in Electric Fields
The sensitivity of Rydberg atoms to electric fields arises due to the high polarizability of Rydberg states. In the presence of electric field, the energy levels shift for different states in a way similar to the splitting of hydrogen atoms. This is called the Stark effect and the associated Stark shifts can be calculated by direct diagonalization of the Hamiltonian matrix. Fig. 2 shows the Stark map of Cs atoms in the vicinity of the $30D$ state. The red dashed line shows the energy level in the absence of electric field. The low-$l$ ($l=0,1,2$) Rydberg states are non-degenerate and separated from the associated hydrogenic manifold. At low electric fields $\mathcal{E}$, the Stark shift is a second-order energy shift and scales quadratically as $\Delta E_{\text{DC}}=\alpha\mathcal{E}^{2}/2$, where $\alpha$ is the DC polarizability of the Rydberg state. For $l>3$, the states are degenerate and therefore experience a linear Stark shift similar to the hydrogen Stark spectrum.
When the electric field interacting with atoms is constant or fluctuates at a low frequency, it is called DC Stark shift. When a high-frequency electric field (RF field) interacts with the atoms, it results in AC Stark shift, which has the same quadratic scaling with the field amplitude $\mathcal{E}$. The magnitude and sign of the AC Stark shift depends on the detuning $\Delta$ of the RF field from the resonant transition frequency and scales as $\Delta E_{\text{AC}}\propto 1/\Delta$. This implies that the magnitude of the shift is larger when the RF electric field is closer to resonance, while its sign is positive (negative) for blue-detuned (red-detuned) RF fields. The Stark effect causes coupling between closely-spaced energy levels leading to the mixing of energy eigenstates.
II-B Principles, Structures and Theories of RAQRs
II-B1 Rydberg-Electromagnetically Induced Transparency (EIT) Spectroscopy and the Standard Scheme[3]
EIT is a quantum interference phenomenon, in which two excitation pathways of a three-level system interfere to open a transmission window in the absorption spectrum of atomic resonance. This is typically accomplished by counter-propagating a pair of laser beams, each of which corresponds to an excitation pathway. Since this is a coherent process, it is ultra-sensitive to phase perturbations and energy level shifts of the three-level system.
Rydberg excitation is prepared via a two-photon excitation process, where a pair of counter-propagating laser beams of different wavelengths, namely the ‘probe’ and ‘coupling (control)’, are overlapped inside an alkali vapor cell, as shown in Fig. 3(a). This is accomplished via EIT in a ladder configuration (see Fig. 3(b)). The probe beam excites the atom from its ground state to an intermediate state, while the coupling beam takes it to a high-energy (principle number $n$) state, namely Rydberg state. When an external RF signal’s frequency is (near) resonant with two adjacent Rydberg states, changes can be observed in the optical spectroscopic signal. As the RF field strength increases from low to high, the Rydberg-EIT transparency peak becomes lower (due to small absorption) and then spilts into two peaks. This is the Autler-Townes splitting (ATS) phenomenon. The corresponding changes observed in the probe signal induced by the addition of a coupling laser and RF field, have been illustrated in Fig. 3(c).
This is the standard scheme (structure) used in RAQR. The electron transitions take place among the four energy levels (quantum states) |$\ket{1}$, |$\ket{2}$, |$\ket{3}$, and |$\ket{4}$ with resonant transition frequencies $\omega_{(P,C,s)}$ and detunings $\Delta_{(P,C)}$, as shown in Fig. 3(b). Each quantum state has its own decay rate, denoted by $\gamma$. The transition |$\ket{1}\rightarrow\ket{2}$ is typically chosen to be D2 transition of Cs (Rb). |$\ket{3}$ could be any Rydberg state, and the excitation is created by tuning the coupling laser wavelength to the appropriate value. Finally, |$\ket{4}$ is the neighboring Rydberg state and its choice is governed by the particular RF frequency that the RAQR would be set up to receive. In this model, the measurement sensitivity approaches the photon shot noise limit, which is $\sim$ µV/cm/$\sqrt{\text{Hz}}$ [3].
II-B2 Superheterodyne Scheme for Enhanced Sensitivity [5, 6, 7, 8]
RAQR’s measurement sensitivity can be further improved by exploiting the superheterodyne principle, where an auxiliary RF signal, namely the local oscillator (LO), is added to the structure to detect relatively weak RF signals. This scheme is presented in Fig. 3(d). The RF signal is typically dozens or hundreds of kHz away from the strong LO in this configuration. The LO facilitates the creating of microwave-dressed Rydberg states and in the presence of two RF signals the frequency splitting measurement is converted into a modulated amplitude measurement, since the Rydberg atoms act as an RF mixer. The LO and the desired RF signal having frequencies $\omega_{(L,s)}$ and detunings of $\Delta_{(L)}$ as well as $\delta_{s}$ (with respect to LO) respectively are shown in Fig. 3(e). The sensitivity has been shown to be in the range of 12.5-55 nV/cm/$\sqrt{\text{Hz}}$ [6]. This method improves the sensitivity and additionally provides access to both amplitude and phase sensing. The phase of the RF signal can be extracted from the output modulated signal, since the phase of LO is known. Similarly, the angle of arrival (AoA) of an RF signal [7] can be accurately determined. When this scheme is extended to multilevel configuration, it allows weak RF signal detection over a wide frequency range of $>$ 1 GHz [8].
In both of the schemes above, a balanced photodetector (at the position of ‘Photodetector’ in Fig. 3) is commonly used to eliminate the laser intensity noise and any noise coming from the atoms in the cell. In the superheterodyne structure, the signal gleaned from balanced photodetection is sent to the lock-in amplifier of Fig. 3(d) for demodulation. Further signal processing may then be performed for accurately inferring the RF amplitude, frequency, phase, and so on.
Figure 3: (a) Standard model of RAQRs and its (b) energy level diagram. (c) Probe laser transmission as a function of probe detuning for three cases - no coupling laser, coupling laser present, both coupling laser and RF field present. (d) Superheterodyne model of RAQRs and its (e) the energy level diagram. High dipole moment of Rydberg atoms is symbolically highlighted in (a) and (d).
II-B3 Theories of RAQRs
The four-level atomic system and its response to electric field can be theoretically simulated by solving Lindblad master equation [6]. The steady state solutions provide the long-term behavior of the system when it is interacting with the environment. The macroscopic quantities of interest can be obtained from these solutions. The different properties of Rydberg atoms may then be studied with the help of the Alkali Rydberg Calculator package[9].
II-C Sources of Impairments
The performance of RAQRs in communication and sensing is influenced by several impairments, as detailed below.
II-C1 RAQR Noise Sources
There are several factors that contribute to the noise of RAQRs. The photon shot noise and the electrical (Johnson-Nyquist) noise from the photodetector may increase the noise floor, making it difficult to detect weak signals. The laser noise plays a crucial role in determining the sensitivity of RAQR. The frequency noise of lasers can be minimized by frequency stabilization techniques, where the laser is locked to the mode of a high-finesse ultra-low expansion (ULE) cavity. The intensity noise of lasers can be reduced similarly by implementing a servo feedback loop and a balanced photodetection scheme. The vibrational instability of mirrors may contribute to the mechanical noise. All these can be mitigated by innovative engineering techniques.
II-C2 Transit Time and Doppler Broadening
The atoms in the vapor cell only interact with laser light during the so called transit time. After this they leave the interaction region due to their thermal motion (kinetic energy) at room temperature. This leads to the broadening of spectral lines termed as transit time broadening.
An atom moving in (against) the direction of light beam will experience an energy level shift i.e., redshifted (blueshifted) with respect to the light frequency. When this frequency is resonant to an atomic transition, light gets absorbed by the atoms obeying the same (Gaussian) velocity distribution of the atoms. These different velocities result in the broadening of spectral lines, termed as Doppler broadening. In RAQR, there is a mismatch in wavelengths for the counter-propagating probe and coupling laser beams which accounts for the Doppler-broadened Rydberg-EIT peak (shown in Fig 3(c)). The broadening is proportional to the residual Doppler mismatch and it is a few MHz.
II-C3 Quantum Projection Noise [3]
It is also known as the quantum shot noise and often referred to as Standard Quantum Limit (SQL), since it sets a lower limit to the RF signal strength that the the RARQ can receive. This noise arises from the quantum fluctuations of the measurement process. The SQL scales as $\propto 1/\sqrt{NT_{r}T_{i}}$, where $N$ is the number of uncorrelated Rydberg atoms participating in RF reception, $T_{i}$ is the integration time and $T_{r}$ is the coherence time of the RAQR. The typical SQL sensitivity is in the range of a few hundred pV/cm/$\sqrt{\text{Hz}}$.
In order to break through the SQL, one has to harness quantum-enhanced measurement systems, such as the complex quantum states of light (Schrodinger cat state, squeezed state, etc) and interacting many-body states. A sensitivity enhancement has been shown recently for a many-body, non-equilibrium Rydberg atomic system [10].
II-C4 Collisional Effects [3]
The interaction or collision of Rydberg atoms with the vapor-cell wall becomes significant at roughly micron-scale distances, since there are boundary conditions imposed on the atomic dipoles, as they approach the wall. There are other collisions in this system, like Rydberg atom - Rydberg atom collisions and ground-state - Rydberg atom collisions, that play an important role. All of these collisions cause dephasing of the quantum states and reduce the coherence time $T_{r}$ of the RAQR receiver. It can be shown that when all these dephasing effects are taken into account, the coherence time $T_{r}$ and hence the sensitivity reaches the maximum value for Rydberg states in the range of $n=45,\cdots,60$.
II-C5 Noise from RF Background
RAQRs cannot differentiate between an RF target signal and the background electromagnetic noise. The change in the response of RAQR in the presence of bandwidth-limited white Gaussian RF noise has been theoretically modelled and verified experimentally. Additionally, by combining it with machine learning models, the RAQR can be trained to receive a particular RF signal of given bandwidth even if the output spectroscopic signal is heavily distorted by noise.
II-D Comparison to Conventional RF Receivers
II-D1 Different Operating Principle
Conventional RF receivers employ antennas for capturing the incident EM waves and to generate currents or voltages. By contrast, the RAQRs realize RF reception via Rydberg excitation created by driving laser beams in an atom vapor cell.
II-D2 SI-Traceable Receiver With Self-Calibration
Conventional antennas need fine calibration for high-integrity reception. The calibrations are performed by placing uncalibrated antennas in a known EM environment that is measured via a well-calibrated antenna, leading to a ‘chicken-or-egg’ dilemma. However, RAQRs allow precise measurements of RF signals which is directly related to Planck’s constant and hence can self-calibrate by relying on SI-traceable atomic standards.
II-D3 Direct Passband to Baseband Conversion
Current RF receivers require tightly-integrated RF front-end components (e.g., mixers, amplifiers, and filters, etc.) for down-converting the passband RF signal to baseband signals, requiring complex integrated circuits, especially for large-scale systems. By contrast, the measurement scheme of RAQRs allow direct down-conversion of RF signals, bypassing the complex integrated circuitry, thus facilitating a simple, compact structure.
II-D4 Different Receiver Impairments
Traditional RF receivers suffer from Gaussian thermal noise and hardware impairments, while RAQRs suffer from a multitude of different impairments - ranging from the noise of lasers and photodetectors to noise arising from broadening and dephasing effects in atomic systems, as explained in Section II-C. The sensitivity for RAQRs is set by the quantum shot noise limit (SQL), whereas for conventional RF receivers the sensitivity is determined by the thermal noise limit. Ref. [11] has argued that the quantum shot noise will duly drop below the thermal noise floor of the conventional RF receivers, when the symbol rate is lower than 1-2 Msymbol/s, provided that all other experimental noise sources are absent.
III RAQR Enabled Quantum-Aided Wireless Communication and Sensing
The application of RAQRs to classical wireless communication and sensing is promising, as indicated by representative studies from diverse research communities that promote quantum-aided classical wireless communication and sensing.
III-A Verified Capabilities
  • C1: Amplitude Detection. RAQRs are capable of measuring the amplitude of an RF signal using the standard scheme discussed in Section II-B1.
  • C2: Phase Detection. It can also retrieve the phase of an RF signal using the superheterodyne principle of Section II-B2.
  • C3: Polarization Detection. The polarization of an RF signal can be detected from the spectroscopic EIT signal, which depends on the angular difference between the laser polarization and the RF wave vector [12]. Moreover, an atomic system relying on polarization selectivity becomes feasible with the aid of metallic structures.
  • C4: Modulation Detection. The amplitude and phase detection capabilities of RAQRs allow the detection of various amplitude/frequency modulation (AM/FM) schemes [13, 12], phase-shift keying (PSK) (BPSK and QPSK [14, 8]), and of quadrature amplitude modulation (QAM) [14].
  • C5: Multiband/Continuous-Band Detection. Simultaneously detecting RF signals in multiple bands becomes feasible using a single RAQR. This can be realized by using different Rydberg final states, where each final state is coupled with a distinct RF band. Alternatively, populating a vapor cell with different atoms and matching each atom species with different laser wavelengths can sense different RF bands [13]. The multiple bands can span over a continuous frequency range through adjacent Rydberg resonance tuning [8].
  • C6: Spatial Displacement and Direction Detection. The spatial displacement of a moving target or angle-of-arrival (AoA) of an incident RF signal[7]) can be detected by RAQRs. By recording the phase-change of the RF signal caused by target movements, one can deduce spatial displacement [15]. and by acquiring the phase-difference of the RF signal between two locations, the AoA can be estimated [7].
TABLE I: State-of-the-art in RAQRs.
                                                              Theoretical studies from the physics community
   Ref.
EIT model
Sensitivity
Atomic
motions
Atomic
collisions
Number of atoms
Vapor cell influence
             Main Contributions
[3]
$\times$
$\surd$
$\times$
$\surd$
$\surd$
$\surd$
(dielectric constant)
Using a bright resonance prepared within an EIT window, it is possible to achieve
high sensitivities
$\sim$
1 µV/cm/$\sqrt{\text{Hz}}$ and detect small RF electric fields
1 µV/cm.
                                                              Experimental studies from the physics community
   Ref.
Structure
Capability
Features
Atom
Energy levels
Laser power
(probe/coupling)
Laser wavelength
(probe/coupling)
RF frequency
Main Contributions
[13]
Standard
C1, C4 (AM, FM),
C5 (multiband)
F1, F2,
F3, F5
133 Cs
85 Rb
(6S 1/2 →6P 3/2
34D 5/2 →35P 3/2 ),
(5S 1/2 →5P 3/2
47D 5/2 →48P 3/2 )
(41.2 µW /
48.7 mW),
(22.3 µW /
43.8 mW)
(850.53 nm /
511.148 nm),
(780.24 nm /
480.27 nm)
(19.626 GHz),
(20.644 GHz)
A multiband Rydberg atomic receiver is
achieved using the combination of two
different atomic species.
[12]
Standard
C1, C3, C4 (AM,
FM, pulsed RF)
F1, F2,
F4
133 Cs
6S 1/2 →6P 3/2
nD 5/2 →(n+1)P 3/2
N/A
852 nm / 510 nm
12.605 GHz,
2.5 GHz
A portable & commercial Rydberg atomic
RF electric-field measurement instrument
is first reported.
[14]
Superheterodyne
C1, C2, C4
(BPSK, QPSK,
QAM)
F1, F3
133 Cs
6S 1/2 →6P 3/2
34D 5/2 →35P 3/2
41.2 µW /
48.7 mW
850.53 nm /
511.148 nm
LO: 19.629 GHz,
IF: 500 kHz
& 1 MHz
The BPSK, QPSK, and QAM modulated
signals are able to be received by RAQRs.
The instantaneous bandwidth of Rydberg
atoms is about 1 MHz - 10 MHz.
[6]
Superheterodyne
C1, C2
F1
133 Cs
6S 1/2 →6P 3/2
47D 5/2 →48P 3/2
120 µW / 34 mW
852 nm / 510 nm
LO: 6.94 GHz,
IF: 150 kHz
Experimentally illustrate a sensitivity of
55 nV/cm/$\sqrt{\text{Hz}}$ and show the minimum
detectable field of 780 pV/cm.
[7]
Superheterodyne
C1, C2, C6 (AoA)
F1, F5
133 Cs
6S 1/2 →6P 3/2
59D 5/2 →59P 3/2
96 µW / 60 mW
852.35 nm /
509.26 nm
19.18 GHz
Superheterodyne Rydberg atomic sensor
is experimentally verified to detect AoA.
[8]
Superheterodyne
C1, C2, C4
(QPSK), C5
(continuous-band)
F1, F2,
F5
87 Rb
5S 1/2 →5P 3/2
70D 5/2 →71P 3/2
→71S 1/2
3.8 µW / 60.4 mW
780 nm / 480 nm
100 MHz -
1.4 GHz
Detect over 1 GHz continuous-band with
best sensitivity of 1.5 µV/cm/$\sqrt{\text{Hz}}$ and
80 dB linear dynamic range.
                                                              Applications to wireless communication and sensing from the communication community
[15]
Superheterodyne
C1, C2, C6
(spatial
displacement)
F1, F2
133 Cs
6S 1/2 →6P 3/2
(66D 5/2 →67P 3/2 ),
(52D 5/2 →53P 3/2 ),
(61D 5/2 →63P 3/2 )
N/A
852 nm / 510 nm
(2.4 GHz),
(5.0 GHz),
(28 GHz)
Improve sensing granularity by an order
of magnitude for WIFI signals and the
28 GHz millimeter wave.
III-B Distinctive Features
  • F1: Extremely-High Sensitivity. Compared to conventional RF receivers, RAQRs exhibit ultra-high sensitivity in detecting RF signals [3, 6], as discussed in Sections II-B1 and II-B2.
  • F2: Broadband Tunability. RAQRs allow a multitude of options in choosing different atomic energy levels combined with different electron transitions to realize the detection of RF signals across a broad frequency range using a single receiver [12, 2].
  • F3: Narrowband Selectivity. The atomic response has a narrowband selectivity, where a narrow instantaneous bandwidth ($\leq 10$ MHz) around the carrier frequency can be realized for coupling between adjacent Rydberg states [14]. This facilitates EM interference-free receptions.
  • F4: Ultra-Wide Input Power Range. RAQRs are capable of receiving not only very weak RF signals, but also very strong RF signals up to several kV/m [12], thereby expanding the linear dynamic range. This ensures the receiver’s operation in high-intensity RF environments.
  • F5: Ultra-High Scalability. RAQRs exhibit an ultra-high scalability and an ability to form advanced receivers (e.g., multiband or continuous-band receivers and arrays) using a single vapor cell. A multiband/continuous-band receiver architecture is facilitated by methods stated in C5 [13, 8]. Construction of a receive array is feasible via splitting laser beams into multiple parallel rays and arranging for each probe-coupling beam pair to propagate through a unique location of the cross-section of a vapor cell [7]. Additionally, the vapor cell can be integrated with metallic structures or embedded into antennas to form a multi-functional receiver (e.g., amplitude-phase-polarization detection) [12].
III-C State-of-the-Art in RAQRs
Next, we review several representative works (see TABLE I), spanning from theoretical aspects to prototype experiments, to applications in communication and sensing, from both the physics and communication communities.
III-C1 Theoretical Studies From the Physics Community [3]
The theoretical studies on RAQRs are mainly from the physics community. A salient aspect of these studies is establishing the Rydberg-EIT signal model by considering a specific 4-level ladder scheme. Another aspect is unveiling the fundamental limitations in approaching the SQL and exploring the optimum sensitivity. Last but not least, several practical limitations must also be considered, such as the noise from lasers and photodetectors, broadening, dephasinng and collisional effects, the influences of vapor cell geometry, etc.
III-C2 Experimental Studies From the Physics Community [14, 6, 13, 12, 7, 8]
Another prevalent branch in the physics community is experimentally designing novel atomic structures to support new capabilities, demonstrate specific features, explore new sensitivity limits, and so forth. The advances of RAQRs have confirmed the feasibility of realizing powerful RF receivers, which will pave the way for more complex upper-layer communication and sensing applications.
III-C3 Applications to Wireless Communication and Sensing From the Communication Community [15]
Attracted by the appealing advantages and inspired by recent advances of RAQRs, upper-layer applications in wireless communication and sensing started to emerge. The authors of [15] demonstrated its superiority in sensing moving objects.
Figure 4: Superheterodyne RAQR based wireless communication transceiver.
IV Rydberg Atomic Quantum SISO and MIMO Schemes
In this section, we introduce the RAQRs into classical wireless systems by detailing the transceiver structure of the RAQ-SISO system, and presenting a potential RAQ-MIMO scheme, which evolved from the schemes portrayed from a physics perspective in [2, 7].
IV-A RAQ-SISO Architecture
Fig. 4 shows each stage of the holistic RAQ-SISO system from a communication perspective. At the transmit side, a classic wireless transmitter converts the baseband signal to a passband RF signal by the following two blocks:
  1. TXB1
    Interpolation: Following Nyquist sampling, the transmitted baseband signal (assuming a bandwidth of $W$) is transformed to its continuous-time counterpart using the sinc function.
  2. TXB2
    Up-conversion: The continuous-time baseband signal of zero center frequency is then upshifted to a passband RF signal of carrier frequency $f_{c}$.
As indicated by the downwards dotted arrows of Fig. 4, the details of the two transmit blocks are seen in the dotted boxes, respectively, where the transmitter relies on an in-phase and quadrature-phase (IQ) modulation scheme, and mainly operates in the digital and RF domains.
The passband RF signal propagates through the wireless channel characterized by a linear system response $h(\tau,t)$, and it is detected by the RAQR based wireless receiver including:
  1. RXB1
    Rydberg Atomic sensing: It realizes RF reception using the superheterodyne RAQR that outputs an optical signal, which can be approximated by an amplified cosine wave of frequency $f_{c}-f_{l}$ [6].
  2. RXB2
    Photodetection: The optical signal is then detected by a photodetector, which can be a direct incoherent optical detection (DIOD) scheme or a balanced coherent optical detection (BCOD) scheme. Both DIOD and BCOD generate an output that can be approximated by a cosine wave of frequency $f_{c}-f_{l}$.
  3. RXB3
    Down-conversion: To obtain a continuous-time baseband signal of the photodetector’s output, down-conversion is realized by a lock-in amplifier, which requires IQ mixers having a mixing frequency of $f_{c}-f_{l}$ and lowpass filters having a bandwidth of $W$.
  4. RXB4
    Sampling: The continuous-time baseband signal is then sampled and quantized by analog-to-digital converters (ADCs) to acquire the baseband samples. The sampling rate obeys Nyquist’s sampling theorem.
  5. RXB5
    Baseband recovery: The baseband samples embody both the intensity and phase of the desired RF signal, which are ultilized for recovering the transmitted baseband signal using specific signal processing algorithms.
The receiver blocks are detailed in their dotted boxes, respectively, where the RAQR based wireless receiver operates in the optical, RF and digital domains.
This architecture is capable of integrating RAQRs into a holistic classical communication system, and paving the way for constructing the corresponding mathematical models and signal processing algorithms.
Figure 5: RAQ-MIMO receiver.
IV-B RAQ-MIMO Architecture
To extend the RAQ-SISO of Fig. 4 to a MIMO architecture, the most straightforward way is to use multiple vapor cells to form an array, where each vapor cell is penetrated using a probe-coupling beam pair to form the Rydberg states for RF reception, as seen in Fig. 4. The optical outputs of all vapor cells are forwarded to the corresponding photodetector, lock-in amplifier and ADC. The outputs of all branches are then appropriately combined to recover the desired RF signal, as seen in Fig. 5(a).
In addition to the above MIMO extension, we introduce a more compact one, as presented in Fig. 5(b). This architecture is implemented using a single vapor cell penetrated by a laser beam array. The optical array signals experience diversity gains owing to their different locations, and are detected by a photodetector array. The outputs are then delivered to multiple lock-in amplifiers and ADCs for baseband recovery. This architecture of the RAQ-MIMO receivers is initially verified in [7] to realize angle-of-arrival detection. This scheme is expected to play an important role in future quantum-aided wireless MIMO systems given its compact form factor.
V Open Problems and Future Directions
Even though significant advances have been achieved in the evolution of RAQRs, the answers to several open problems remain to be further explored. We thus list a number of future directions of significant interest as follows. Advanced designs of RAQR have the potential of approaching or even exceeding the SQL, thereby outperforming the Shannon-limit of the current RF receivers.
Theoretical Modeling of RAQR-Aided Wireless Systems: At the time of writing, the integration of RAQRs into communication and sensing systems is in its infancy. The theoretical modelling is mainly dedicated to the input-output relationship of the vapor cell. However, a holistic baseband equivalent mathematical model, especially the baseband equivalent model, has to be developed for the RF signals of various systems (e.g., SISO, MIMO).
Signal Processing of RAQR-Aided Wireless Systems: Furthermore, computationally efficient signal processing algorithms must be conceived for communication-/sensing-oriented functionalities, including signal detection, channel estimation, and spatial displacement/direction detection etc.
Experimental Implementations and Validations: Previous efforts have shown the potential of RAQRs, but substantive future efforts are required for experimentally investigating its practical capability. Furthermore, the attainable system-level benefits must be quantified.
Cost, Size, Weight, and Power Reduction: To commercialize RAQRs, mature cost-saving, miniaturization and power-conserving technologies of the laser generators and their accessories are required to realize such receivers.
VI Conclusions
RAQRs and their integration into wireless communication and sensing were considered. We briefly introduced the fundamentals of Rydberg atoms and discussed the RAQRs’ capabilities as well as distinctive features, followed by a contemporary survey of the state-of-the-art studies. To support the integration of RAQRs with classical wireless systems, we advocated RAQ-SISO and RAQ-MIMO schemes. Finally, we concluded by proposing promising research directions.